OpenStudy (xbeenchantedx):

how do you solve by completing the sq this problem: 4x^2 + 12x + 19 = 11

OpenStudy (anonymous):

first subtract 19, then divide by 4

OpenStudy (anonymous):

\[4x^2 + 12x + 19 = 11\] \[4x^2 + 12x = -8\] \[x^2+3x=-2\]

OpenStudy (anonymous):

then complete the square \[(x+\frac{3}{2})^2=-2+\frac{9}{4}\] \[(x+\frac{3}{2})^2=\frac{1}{4}\]

OpenStudy (anonymous):

take the square root and get \[x+\frac{3}{2}=\pm\frac{1}{2}\] then solve the two equation \[x+\frac{3}{2}=\frac{1}{2}\] and \[x+\frac{3}{2}=-\frac{1}{2}\] and you are done